# Towards Explaining Deep Neural Network Compression Through a Probabilistic Latent Space

### v1

@@ 5.3 Multivariate T-Student Distribution
Figure 4: Comparing AP3 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
5.4 Comparing AP2 and AP3
Our goal is to compare AP2 and AP3 metrics to distinguish lowest, highest, and random pruning methods. Both metrics share the same covariance matrix for a fair comparison. We evaluate identical pruning percentages side by side using AP2 and AP3. Consequently, by leveraging AP2, we can gain a more comprehensive insight into which pruning method exhibits superior performance differences. The same comparison for pruned AlexNet is shown in Appendix.
@@ 6 Conclusion
In this study, we investigated network sparsity interpretability by employing probabilistic latent spaces. Our main aim was to explore the concept of network pruning using the Kullback-Leibler (KL) divergence of two probabilistic distributions associated with the original network and its pruned counterpart, utilizing their weight tensors as parameters for the latent spaces. Throughout a series of experiments, we demonstrated the efficacy of two projective patterns, namely AP2 and AP3, in elucidating the sparsity of weight matrices. Our analysis provided valuable explanatory insights into the distributions of pruned and unpruned networks. Remarkably, our results indicated that, despite certain limitations encountered in our experiments compared to theoretical claims, nearly all three pruning methods we employed exhibited alignment with the expectations suggested by Lemma 2 when we use the nearly optimal sparsity level.
@@
Going ahead, we intend to thoroughly investigate the implications of optimal pruning for preserving strong information flow  [27, 28, 29] between layers through the use of probabilistic latent spaces. Our main goal is to substantiate the claim that optimal pruning ensures the retention of highly comparable mutual information between the layers in the pruned and original networks. By this, we intend to use our findings in this work to better tackle the information transfer challenges in domain adaptation problems.
@@
Figure 5: Comparing AP3 and AP2 (Y-axis shows both values) for pruned VGG16 on CIFAR10.
7 Acknowledgement

### v3

@@ 5.3. Multivariate T-Student Distribution
Figure 4. Comparing AP2 & PD for pruned ResNet50 and VGG16 on Tiny-ImageNet. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on VGG16 (Left) and ResNet50(Right).
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5. Comparing AP3 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
Figure 6. Comparing AP3 & PD for pruned ViT on Tiny-ImageNet. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning.
5.4. Comparing AP2 and AP3
Our goal is to compare AP2 and AP3 metrics to distinguish the lowest, highest, and random pruning methods. Both metrics share the same covariance matrix for a fair comparison. We evaluate identical pruning percentages side by side using AP2 and AP3. Consequently, by leveraging AP2, we can gain a more comprehensive insight into which pruning method exhibits superior performance differences. The same comparison for pruned AlexNet is shown in the Appendix.
@@ 6. Conclusion
In this study, we investigated network sparsity interpretability by employing probabilistic latent spaces. Our main aim was to explore the concept of network pruning using the Kullback-Leibler (KL) divergence of two probabilistic distributions associated with the original network and its pruned counterpart, utilizing their weight tensors as parameters for the latent spaces. Throughout a series of experiments, we demonstrated the efficacy of two projective patterns, namely AP2 and AP3, in elucidating the sparsity of weight matrices. Our analysis provided valuable explanatory insights into the distributions of pruned and unpruned networks. Remarkably, our results indicated that, despite certain limitations encountered in our experiments compared to theoretical claims, nearly all three pruning methods we employed exhibited alignment with the expectations suggested by Lemma 3.1 when we use the nearly optimal sparsity level.
@@
Going ahead, we intend to thoroughly investigate the implications of optimal pruning for preserving strong information flow  (Ganesh et al., 2022, 2021; Andle and Yasaei Sekeh, 2022) between layers through the use of probabilistic latent spaces. Our main goal is to substantiate the claim that optimal pruning ensures the retention of highly comparable mutual information between the layers in the pruned and original networks. By this, we intend to use our findings in this work to better tackle the information transfer challenges in domain adaptation problems.
@@
Figure 7. Comparing AP3 and AP2 (Y-axis shows both values) for pruned VGG16 on CIFAR10.
Acknowledgments